OpenStudy (anonymous):

A number less it reciprocal is 8/3. What is the number?

Directrix (directrix):

Let n be the number. The reciprocal of n is 1/n provided n is not zero. So, solve this equation for n: n - (1/n) = (8/3)

OpenStudy (anonymous):

n-(1/n)-(8/3)=0 I can't factor it so what will i do next?

Directrix (directrix):

Take the equation: n - (1/n) = (8/3) --> multiply each member by (3n) That will clear the fractions. Post what you get, okay?

Directrix (directrix):

@danielle02 How is the equation solving coming along?

OpenStudy (anonymous):

3x^2-8x-3=0

OpenStudy (anonymous):

@Directrix is that correct?

Directrix (directrix):

Yes. N is now x. Next step: Factor 3x^2-8x-3=0

Directrix (directrix):

What are two numbers that multiply to -9 and add to -8?

OpenStudy (anonymous):

(x-3)(3x+1)

OpenStudy (anonymous):

x1=3 and x2=-1/3

Directrix (directrix):

I think there will be two sets of solutions. Let's see.

OpenStudy (anonymous):

If u need help we will explain

OpenStudy (anonymous):

Thanks!

Directrix (directrix):

>> A number less it reciprocal is 8/3. If x = 3, then 3 -1/3 has to be 8/3. If x = -1/3, then -1/3 - 3/1 has to be 8/3.

Directrix (directrix):

Now, I think there is only one solution because this is not true: If x = -1/3, then -1/3 - 3/1 has to be 8/3.

Directrix (directrix):

So, the number would just be 3.